OpenStudy (anonymous):

2. A manufacturer of electronic equipment requires a certain rare metal. He has a reserve supply of 4,000,000 units that he will not be able to replace. If the rate at which the metal is used is given by f(t)=100,000e.03t where t is time in years, how long will it be before he uses his supply?

OpenStudy (anonymous):

Set up an inequality: \[4000000 \ge 100000e^(3t)\]

OpenStudy (anonymous):

Remember you can dispel of e^x by taking the natural log.

OpenStudy (anonymous):

\[\ln(e^x) = x\]